### v1

# Synthesizing Grid Data with Cyber Resilience and Privacy Guarantees

Abstract
Differential privacy (DP) provides a principled approach to synthesizing data (e.g., loads) from real-world power systems while limiting the exposure of sensitive information. However, adversaries may exploit synthetic data to calibrate cyberattacks on the source grids. To control these risks, we propose new DP algorithms for synthesizing data that provide the source grids with both cyber resilience and privacy guarantees. The algorithms incorporate both normal operation and attack optimization models to balance the fidelity of synthesized data and cyber resilience. The resulting post-processing optimization is reformulated as a robust optimization problem, which is compatible with the exponential mechanism of DP to moderate its computational burden.
I Introduction
Optimal power flow (OPF) analysis in power systems requires realistic grid models with accurate network, generation, and load parameters—data that is extremely difficult to source from real-world grids due to privacy and cybersecurity concerns. While the lack of such models has inspired the development of artificial grids [1, 2, 3], a more principled approach leverages the theory of differential privacy (DP) [4] to release grid models directly from real-world systems.
The DP theory asserts that it is impossible—up to prescribed privacy parameters—to infer the original parameters from their DP release. Such strong privacy guarantees originate from Laplacian perturbations [5] of real grid parameters, followed by post-processing optimization of the perturbed parameters to restore their modeling fidelity to the source grid, e.g., in terms of similarity of the OPF outcomes [6, 7, 8]. The DP theory also lies at the core of modern privacy-preserving OPF solvers [9, 10, 11], the release of aggregated grid statistics [12], and related grid information [13].
However, the privacy guarantees alone may not suffice to release grid parameters, as cybersecurity risks associated with such releases remain largely unexplored. Possible cyber attacks include false data injection, which subtly alters state estimation results [14], and related load redistribution, which manipulates demand measurements to increase OPF cost and constraint violation [15]. While executing such attacks requires some grid knowledge [16], traditionally, this information has been difficult to obtain. However, the availability of synthetic grid parameters may unintentionally inform adversaries and help in calibrating the attack.
Contribution: Recognizing the risks that synthetic grid parameters may inform cyber adversaries, we develop new DP algorithms that simultaneously guarantee cyber resilience and privacy for the source power grids. Our algorithms build on [6, 7, 8] and leverage the Laplace mechanism and post-processing optimization to tune synthetic data while anticipating cyber risks through embedded attack optimization.
The contributions of this paper are summarized as follows:
  1. 1.
    We formulate a Cyber Resilient Obfuscation (CRO) algorithm, an optimization-based algorithm to release electric load data with a guarantee to preserve the privacy of the original data and ensure the resilience of the source grid to load redistribution attacks. The algorithm post-processes synthetic loads by balancing their fidelity with the potential damage to the grid.
  2. 2.
    The underlying post-processing optimization is an intractable trilevel problem, which is reduced to a tractable yet more conservative bilevel problem. We achieve this by exploring the connections between robust and bilevel optimization, in the spirit of [17].
  3. 3.
    To further improve computational tractability of the algorithm, we provide the extension of CRO, termed CRO-Exp, which uses the exponential mechanism of DP to identify only the most important constraints for post-processing optimization of synthetic loads.
We next provide preliminaries on OPF and DP theory. Sec. III explains the risks of cyberattacks, and Sec. IV introduces the algorithms to mitigate them. Sec. V presents simulations, and Sec. VI concludes. Proofs are relegated to the appendix.
Notation: lower- (upper-) case boldface letters denote column vectors (matrices). Scalar $a_{i}$ is the $i^{\text{th}}$ element of vector $\mathbf{a}$, while vector $\mathbf{a}_{i}$ collects the $i^{\text{th}}$ row of matrix $\mathbf{A}$. Vectors $\mathbf{0}$ and $\mathbf{1}$ are the all-zero and all-one vectors; $\top$ stands for transposition, and $\mathbf{x}^{\star}$ denotes the optimal value of $\mathbf{x}$.
II Preliminaries
II-A Optimal Power Flow (OPF) Problem
The OPF problem seeks the least-cost generation dispatch that meets electrical loads while satisfying grid constraints:
$C_{\text{opf}}(\mathbf{d})=\underset{{\mathbf{p},\mathbf{v}\geqslant\mathbf{0}}}{\text{minimize}}\quad$ $\mathbf{q}^{\top}\mathbf{p}+\bm{\psi}^{\top}\mathbf{v}$ (1a)
subject to $\overline{\mathbf{p}}\leqslant\mathbf{p}\leqslant\overline{\mathbf{p}}$ (1b)
$\bm{1}^{\top}(\mathbf{p}-\mathbf{d})=0$ (1c)
$|\mathbf{F}(\mathbf{p}-\mathbf{d})|\leqslant\overline{\mathbf{f}}+\mathbf{v}$ (1d)
The model maps the load vector $\mathbf{d}$ to the optimal dispatch cost $C_{\text{opf}}(\mathbf{d})$. The decision variables include the vector $\mathbf{p}$ of generator dispatch and vector $\mathbf{v}$ of power flow constraint violations. Dispatch costs are linear with coefficients $\mathbf{q}$. Constraint (1b) requires generator dispatch $\mathbf{p}$ to remain within a feasible range of $[\underline{\mathbf{p}},\overline{\mathbf{p}}]$. Constraint (1c) defines system-wide power balance between dispatched generation and loads. The power flows in transmission lines are computed using the matrix of power transfer distribution factors $\mathbf{F}$, which maps net power injections $(\mathbf{p}-\mathbf{d})$ to power flows as $\mathbf{F}(\mathbf{p}-\mathbf{d})$. They are capped by line capacity $\overline{\mathbf{f}}$ using constraint (1d). In highly loaded condition, these constraints can be temporally violated by $\mathbf{v}\geqslant\mathbf{0}$. Transmission constraint violations are not desired, so they are penalized in the objective function with a typically large parameter $\bm{\psi}$, i.e., $|\bm{\psi}|\gg|\mathbf{c}|$.
We write the linear OPF problem (1) in a compact form
$C_{\text{opf}}(\mathbf{d})=\underset{{\mathbf{x}}}{\text{minimize}}\quad$ $\mathbf{c}^{\top}\mathbf{x}$ (2a)
subject to $\mathbf{a}_{k}^{\top}\mathbf{x}+\mathbf{b}_{k}^{\top}\mathbf{d}+e_{k}\leqslant\mathbf{0},$
$\quad\quad\quad\!\;\forall k=1,\dots,K,$ (2b)
where $\mathbf{x}=[\mathbf{p}^{\top}\;\mathbf{v}^{\top}]^{\top}$, $\mathbf{c}=[\mathbf{q}^{\top}\;\bm{\psi}^{\top}]^{\top}$. The $K$ inequalities in (2b) encode the dispatch constraints (1b), power flow constraints (1c) and (1d) using properly specified parameters $\mathbf{a}_{1},\mathbf{b}_{1},e_{1},\dots,\mathbf{a}_{K},\mathbf{b}_{K},e_{K}$. We make the following non-restrictive assumption on OPF.
Assumption 1.
For a feasible vector $\mathbf{d}$, problem (2) satisfies the linearly independent constraint qualification (LICQ).
II-B Differential Privacy for Synthetic OPF Datasets
Optimization parameters in problem (1) are either classified or owned by private system actors, and thus can not be directly disclosed to public. Our goal is thus to synthesize some realistic version of these parameters. Without much loss of generality, we focus on the obfuscation of demand vector $\mathbf{d}$, while other parameters can be synthesized similarly. The state-of-the-art obfuscation algorithms synthesize high-quality demand vectors $\tilde{\mathbf{d}}$, which do not disclose the original vector $\mathbf{d}$ yet behave consistently in terms of power flow analysis [8, 6, 7]. They leverage DP to render original and synthetic vectors $\mathbf{d}$ and $\tilde{\mathbf{d}}$ statistically indistinguishable, up to some prescribed non-negative parameters $\alpha$ and $\varepsilon$. The former refers to adjacency of two arbitrary vectors [4].
Definition 1 (Adjacency).
Two vectors $\mathbf{d},\mathbf{d}^{\prime}\in\mathcal{D}\subset\mathbb{R}^{n}$ are $\alpha-$adjacent, if $\exists i\in 1,\dots,n$, such that $d_{j}=d_{j}^{\prime},\forall j\in\{1,\dots,n\}\backslash i$, and $|d_{i}-d_{i}^{\prime}|\leqslant\alpha$ for $\alpha>0$. That is, they are different in one item by at most $\alpha$. $\triangleleft$
To synthesize a DP version $\tilde{\mathbf{d}}$ of $\mathbf{d}$, the standard Laplace mechanism applies a random noise to the original data, i.e., $\tilde{\mathbf{d}}=\mathbf{d}+\text{Lap}(\frac{\alpha}{\varepsilon})^{n}$, where $\text{Lap}(s)^{n}$ is the random draw from the $n-$dimensional Laplace distribution with zero mean and scale $s$ [7]. This mechanism satisfies the definition of $\varepsilon-$DP.
Definition 2 ($\varepsilon-$DP).
A Laplacian mechanism is called $\varepsilon-$DP if, for any outcome $\tilde{\mathbf{d}}$ in the range of the mechanism, and any two $\alpha-$adjacent load vectors $\mathbf{d}$ and $\mathbf{d}^{\prime}$, the ratio of probabilities of observing the same outcome is bounded as
$\frac{\text{Pr}[\mathbf{d}{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}^{\prime}}+\text{Lap}(\frac{\alpha}{\varepsilon})^{n}=\tilde{\mathbf{d}}]}{\text{Pr}[\mathbf{d}^{\prime}+\text{Lap}(\frac{\alpha}{\varepsilon})^{n}=\tilde{\mathbf{d}}]}\leqslant\text{exp}(\varepsilon).$ (3)
where $\varepsilon$ is a prescribed non-negative parameter. $\triangleleft$
Parameter $\varepsilon$ is termed privacy loss. Intuitively, a smaller $\varepsilon$ results in more noise applied to data and higher requirement for distribution similarity, which would make it more likely to observe the same random outcome. However, the Laplace mechanism alone is likely to produce such load vector $\tilde{\mathbf{d}}$ that does not admit a feasible OPF solution, i.e., $C_{\text{opf}}(\tilde{\mathbf{d}})=\varnothing.$ The prior work introduced the following two-stage solution:
  1. 1.
    Laplace mechanism $\tilde{\mathbf{d}}^{0}=\mathbf{d}+\text{Lap}\left(\!\frac{2\alpha}{\varepsilon}\!\right)^{n}$, followed by
  2. 2.
    Post-processing of $\tilde{\mathbf{d}}^{0}$ using a bilevel optimization:
$\underset{{\tilde{\mathbf{d}}}}{\text{minimize}}\;\;$ $\lVert C_{\text{opf}}(\tilde{\mathbf{d}})-\tilde{C}_{\text{opf}}\rVert_{1}+\gamma\lVert\tilde{\mathbf{d}}-\tilde{\mathbf{d}}^{0}\rVert_{1}$ (4a)
where the OPF costs come from
$C_{\text{opf}}(\tilde{\mathbf{d}})=\underset{{\mathbf{x}}}{\text{minimize}}\quad$ $\mathbf{c}^{\top}\mathbf{x}$ (4b)
subject to $\mathbf{a}_{k}^{\top}\mathbf{x}+\mathbf{b}_{k}^{\top}\tilde{\mathbf{d}}+e_{k}\leqslant\mathbf{0},\forall k$ (4c)
The synthetic vector $\tilde{\mathbf{d}}$ is optimized using feedback from the embedded OPF problem, which constraints $\tilde{\mathbf{d}}$ to take only those values that admit a feasible OPF solution. The main objective of (4) is to match the OPF cost on synthetic load vector with that on the original load vector, thereby ensuring high modeling fidelity of the synthetic data. For that purpose, (4) includes a differentially private estimate of OPF costs on true data $\tilde{C}_{\text{opf}}=C_{\text{opf}}(\mathbf{d})+\text{Lap}(\frac{2\overline{c}}{\varepsilon}),$ where $\overline{c}$ is the cost of the most expensive generator. The second term in (4a) is non-dominant and used to regularize the synthetic dataset with some small hyper-parameter $\gamma>0$. Solution to (4) is the feasible and cost-consistent synthetic counterpart $\tilde{\mathbf{d}}$, which ensures $\varepsilon-$DP guarantee for the original load vector $\mathbf{d}$ [8].
One barrier to releasing synthetic OPF parameters is the risk posed by cyber adversaries who might exploit them to disrupt grid operations. Next, we substantiate these risks.
III Cyber Resilience Risks in Releasing Differentially Private OPF Datasets
Although synthetic OPF datasets contribute to overall grid transparency and enable independent power flow analysis, they can also be misused by cyber adversaries launching attacks on the grid. One class of attacks, which is of interest to this work, is load redistribution attacks. In terms of OPF problem (2), the adversary optimizes an attack vector $\bm{\delta}$ that alters loads in $\mathbf{d}$ to increase either the dispatch cost or the magnitude of power flow constraint violations.
According to [15], the optimal attack vector is found by solving the following bilevel optimization (BO) problem:
$C_{\text{att}}^{\text{BO}}(\mathbf{d})=\underset{{\bm{\delta}\in\Delta}}{\text{maximize}}\;C_{\text{opf}}(\mathbf{d}+\bm{\delta})$ (5a)
where the OPF costs come from
$C_{\text{opf}}(\mathbf{d}+\bm{\delta})=\underset{{\mathbf{x}}}{\text{minimize}}\;$ $\mathbf{c}^{\top}\mathbf{x}$ (5b)
subject to $\mathbf{a}_{k}^{\top}\mathbf{x}+\mathbf{b}_{k}^{\top}(\mathbf{d}+\bm{\delta})+e_{k}\leqslant\mathbf{0},\!\!\!$ (5c)
$\forall k.$ The attack vector $\bm{\delta}$ alters nominal loads using feedback from the embedded OPF problem, and is constrained by the set of admissible attacks
$\Delta$$\triangleq\left\{\bm{\delta}\;\bigg{|}\begin{array}[]{l}\underline{\bm{\delta}}\leqslant\bm{\delta}\leqslant\overline{\bm{\delta}}\\ \bm{1}^{\top}\bm{\delta}=0\end{array}\right\}$ (5f)
where $\overline{\bm{\delta}}$ and $\underline{\bm{\delta}}$ are limits on attack magnitude, and $\bm{1}^{\top}\bm{\delta}=0$ ensures that the total system loading remains unchanged after the attack, thus ensuring the stealthiness of the attack.
While the actual load vector $\mathbf{d}$ is not revealed to public, the adversary may leverage its DP release $\tilde{\mathbf{d}}$ to calibrate the attack. Our experiments in Sec. V reveal that the vector computed on $\tilde{\mathbf{d}}$ leads to a substantial increase of OPF costs across standard power systems benchmarks (see Tab. I).
IV Cyber Resilience and Privacy Guarantees for Synthetic OPF Datasets
Recognizing the risks of misusing synthetic datasets, we revisit the post-processing to enhance the cyber resilience of source grids. Instead of (4), we propose the following upper-level objective for the post-processing optimization:
$\underset{{\tilde{\mathbf{d}}}}{\text{minimize}}\;$ $\lVert C_{\text{opf}}(\tilde{\mathbf{d}})-\tilde{C}_{\text{opf}}\rVert_{1}+\beta\lVert C_{\text{att}}^{\text{BO}}(\tilde{\mathbf{d}})-\tilde{C}_{\text{opf}}\rVert_{1}$ (6)
where the first term control the fidelity of the synthetic data, and the second term measures the damage under attack, calibrated on the synthetic data. The objective is thus a trade-off between the fidelity of synthetic grid parameters and resilience of the grid to load redistribution attacks, which can be explored by varying parameter $\beta>0.$ The challenge, however, is that (6) requires solving a trilevel optimization problem, where the synthetic data is optimized over embedded BO attack model $C_{\text{att}}^{\text{BO}}(\tilde{\mathbf{d}})$. Inspired by [17], we seek computational tractability by exploring the connection between the bilevel model of attack and robust optimization.
IV-A Computational Tractability via Robust Optimization (RO)
The conservative RO approximation of (5) is
$C_{\text{att}}^{\text{RO}}(\mathbf{d})=$ $\;\underset{{\mathbf{x}}}{\text{minimize}}\;\;\mathbf{c}^{\top}\mathbf{x}$ (7a)
subject to $\underset{\bm{\delta\in\Delta}}{\max}\left[\mathbf{a}_{k}^{\top}\mathbf{x}+\mathbf{b}_{k}^{\top}(\mathbf{d}+\bm{\delta})+e_{k}\right]\leqslant\mathbf{0},\forall k,\!\!\!$ (7b)
where each constraint $k$ is formulated for the worst-case realization of the attack vector from the set of admissible attacks. In contrast to bilevel formulation (5), the RO attack is specific to each constraint. The following result shows that the RO attack provides an upper-bound on the BO attack.
Proposition 3 (Conservative attack approximation).
For any feasible load vector $\mathbf{d}$, relation $C_{\text{att}}^{\text{RO}}(\mathbf{d})\geqslant C_{\text{att}}^{\text{BO}}(\mathbf{d})$ holds. $\triangleleft$
Although conservative, formulation (7) is computationally advantageous over (5) as it admits a linear programming reformulation via duality [18]. Let $\underline{\bm{\mu}}$ and $\overline{\bm{\mu}}$ be the duals of the first two constraints in (5f), and $\lambda$ be the dual of the last condition in (5f). The exact reformulation of (7) is
$C_{\text{att}}^{\text{RO}}(\mathbf{d})=\;\underset{{\mathbf{x},\underline{\bm{\mu}},\overline{\bm{\mu}},\lambda}}{\text{minimize}}\;\;\;$ $\mathbf{c}^{\top}\mathbf{x}$ (8a)
subject to $\mathbf{a}_{k}^{\top}\mathbf{x}+\mathbf{b}_{k}^{\top}\mathbf{d}$
$\;+\overline{\bm{\mu}}_{k}^{\top}\overline{\bm{\delta}}-\underline{\bm{\mu}}_{k}^{\top}\underline{\bm{\delta}}+e_{k}\leqslant\mathbf{0},\!\!\!$ (8b)
$\mathbf{b}_{k}-\overline{\bm{\mu}}_{k}+\underline{\bm{\mu}}_{k}-\bm{1}\lambda_{k}=\bm{0},$ (8c)
$\underline{\bm{\mu}}_{k},\overline{\bm{\mu}}_{k}\geqslant\mathbf{0},\forall k=1,\dots,K.$ (8d)
Therefore, replacing $C_{\text{att}}^{\text{BO}}$ with $C_{\text{att}}^{\text{RO}}$ in objective function (6) gives rise to bilevel post-processing optimization, which can be handled by mixed-integer optimization solvers [8, 7].
Next, we introduce a tractable post-processing algorithm for synthesizing loads with privacy and cyber resilience guarantees. Then, in Sec. IV-C, we modify the algorithm to tune the computational burden of the RO approximation.
IV-B Differentially Private CRO
Input: $\mathbf{d}$, $(\alpha,\varepsilon_{1},\varepsilon_{2})$, $(\beta,\gamma,\Delta)$
1 Initial load obfuscation: $\tilde{\mathbf{d}}^{0}=\mathbf{d}+\text{Lap}\left(\!\frac{\alpha}{\varepsilon_{1}}\!\right)^{n}$
2 DP estimation of OPF costs: $\tilde{C}_{\text{opf}}=C_{\text{opf}}(\mathbf{d})+\text{Lap}\left(\frac{\alpha\overline{c}}{\varepsilon_{2}}\right)$
3 Post-processing optimization of the synthetic load vector
$\tilde{\mathbf{d}}\in\underset{\tilde{\mathbf{d}}}{\text{argmin}}\;$ $\lVert C_{\text{opf}}(\tilde{\mathbf{d}})-\tilde{C}_{\text{opf}}\rVert_{1}$
$\quad+\beta\lVert C_{\text{att}}^{\text{RO}}(\tilde{\mathbf{d}})-\tilde{C}_{\text{opf}}\rVert_{1}+\gamma\lVert\tilde{\mathbf{d}}-\tilde{\mathbf{d}}^{0}\rVert_{1}$ (9)
Output: Synthetic load vector $\tilde{\mathbf{d}}$
Algorithm 1 Privacy-preserving CRO
The CRO algorithm for privacy-preserving and cyber-resilient synthesis of load parameters is summarized in Alg. 1. It takes as inputs load adjacency and $\varepsilon$-DP parameters, as well as optimization trade-off, regularization and attack parameters, $\beta$, $\gamma$ and $\Delta$, respectively. Step 1 initializes the synthetic load vector using the Laplace mechanism with a privacy loss of $\varepsilon_{1}$. Step 2 performs a DP estimation of the OPF costs on real loads using the Laplace mechanism with a privacy loss of $\varepsilon_{2}$. Following prior work in [8], this step requires the cost $\overline{c}$ of the most expensive generator. Finally, Step 3 post-processes the initial synthetic load by solving the bilevel optimization problem (1) using the conservative RO approximation of the attack. Since Step 3 does not optimize over real data, it does not introduce any privacy loss. The complete formulation of (1) can be seen in Appendix B.
The resilience of the source grid to load redistribution attacks is controlled by the parameter $\beta$ and admissible set $\Delta$. Naturally, a larger $\beta$ and a larger set $\Delta$ lead to greater resilience, but at the expense of the fidelity of the synthesized data. Our experiments in Sec. V will justify for the choices of these parameters. The privacy guarantee for $\alpha$-adjacent load vectors is established by the following result.
Theorem 4 (DP of CRO).
Setting $\varepsilon_{1}=\varepsilon_{2}=\varepsilon/2$ renders Alg. 1 $\varepsilon-$DP for $\alpha-$adjacent load vectors. $\triangleleft$
IV-C Exponential Mechanism to Ease Computational Burden
While the RO approximation (7) leads to a more tractable bilevel optimization, it is still computationally expensive in large systems due to the massive amount of variables and complementarity constraints, as later substantiated by Fig. 2. We propose to alleviate the computational burden by selecting only a subset $\mathcal{K}=\{k_{i}\}_{i=1}^{\tau}$ of $\tau$ constraints for RO reformulation that affect the OPF cost the most. The remaining constraints $\mathcal{K}^{\prime}:=\{1,\dots,K\}\backslash\mathcal{K}$ are enforced deterministically. Setting $\tau=K$ leads to the full RO formulation, while $\tau<K$ leads to a reduced problem:
$C_{\text{att},\tau}^{\text{RO}}(\mathbf{d})=$ $\;\underset{{\mathbf{x}}}{\text{minimize}}\;\;\mathbf{c}^{\top}\mathbf{x}$ (10a)
subject to $\underset{\bm{\delta\in\Delta}}{\max}\left[\mathbf{a}_{k}^{\top}\mathbf{x}+\mathbf{b}_{k}^{\top}(\mathbf{d}+\bm{\delta})+e_{k}\right]\leqslant\mathbf{0},$ (10b)
$\mathbf{a}_{k^{\prime}}^{\top}\mathbf{x}+\mathbf{b}_{k^{\prime}}^{\top}\mathbf{d}+e_{k^{\prime}}\leqslant\mathbf{0},$ (10c)
$\quad\quad\quad\quad\quad\quad\quad\!\;\forall k\in\mathcal{K},\;\forall k^{\prime}\in\mathcal{K^{\prime}}.$
While directly replacing $C_{\text{att}}^{\text{RO}}(\tilde{\mathbf{d}})$ with $C_{\text{att},\tau}^{\text{RO}}(\tilde{\mathbf{d}})$ in Alg. 1 alleviates the computational burden, this also degrades the privacy guarantee of Theorem 4: since the worst-case constraint set $\mathcal{K}$ is specific to a particular load vector $\mathbf{d}$, the post-processing on $\mathcal{K}$ would leak information we intend to obfuscate. As a remedy, we leverage the report-noisy-max algorithm, a discrete version of the exponential mechanism of DP [4], to privately identify the worst-case constraints without leaking information about the actual load. The resulting algorithm, termed CRO-Exp, is given in Alg. 2.
The first two steps of Alg. 2 follow those in Alg. 1. At Step 3, the algorithm applies the exponential mechanism $\tau$ times to construct set $\mathcal{K}$. In each iteration $t$, the mechanism identifies the constraint $k_{t}$ that—when reformulated in a robust fashion—leads to the greatest increase of OPF cost. After $\tau$ iterations, set $\mathcal{K}$ contains $\tau$ worst-case constraints. Finally, Step 4 solves the post-processing optimization with only $\tau$ constraints reformulated in RO way.
Theorem 5 (DP of CRO-Exp).
Setting $\varepsilon_{1}=\varepsilon_{2}=\varepsilon/3$ and $\varepsilon_{3}=\varepsilon/(3\tau)$ renders Alg. 2 $\varepsilon$-DP for $\alpha$-adjacent loads. $\triangleleft$
Input: $\mathbf{d}$, $(\alpha,\varepsilon_{1},\varepsilon_{2},\varepsilon_{3})$, $(\beta,\gamma,\Delta,\tau)$, $\mathcal{K}=\{\varnothing\}$
1 Initial load obfuscation: $\tilde{\mathbf{d}}^{0}=\mathbf{d}+\text{Lap}\left(\!\frac{\alpha}{\varepsilon_{1}}\!\right)^{n}$
2 DP estimation of OPF costs: $\tilde{C}_{\text{opf}}=C_{\text{opf}}(\mathbf{d})+\text{Lap}\left(\frac{\alpha\overline{c}}{\varepsilon_{2}}\right)$
3 DP estimation of the set $\mathcal{K}$ of the worst-case constraints
for $t=1,\dots,\tau$ do
       for $k=1,\dots,K$ do
             $C_{k}=C_{\text{att},t}^{\text{RO}}(\mathbf{d})+\text{Lap}\left(\frac{\alpha\overline{c}}{\varepsilon_{3}}\right)$
       end for
      $k_{t}\leftarrow\text{argmax}_{k}\;C_{k}$
      [2pt] $\mathcal{K}\leftarrow\mathcal{K}\cup\{k_{t}\}$
end for
4 Post-processing optimization of the synthetic load vector
$\tilde{\mathbf{d}}\in\underset{\tilde{\mathbf{d}}}{\text{argmin}}\;$ $\lVert C_{\text{opf}}(\tilde{\mathbf{d}})-\tilde{C}_{\text{opf}}\rVert_{1}$
$\hskip 4.5pt+\beta\lVert C_{\text{att},\tau}^{\text{RO}}(\tilde{\mathbf{d}})-\tilde{C}_{\text{opf}}\rVert_{1}+\gamma\lVert\tilde{\mathbf{d}}-\tilde{\mathbf{d}}^{0}\rVert_{1}$ (11)
Output: Synthetic load vector $\tilde{\mathbf{d}}$
Algorithm 2 Privacy-preserving CRO-Exp
V Experiment Results
We run experiments using standard power grid testbeds. The set of admissible attacks includes the limits on attack magnitude as percentage $\eta$ of nominal loads. The privacy loss $\varepsilon=1$, and we vary adjacency $\alpha$ throughout the experiments. The code and data to replicate our results are available at
https://github.com/Wu-ShengY/CRO_SynDataset.
V-A Substantiating Attacks Calibrated on DP Data
Table I collects the damage of load redistribution attacks on the actual $\mathbf{d}$ and synthetic $\tilde{\mathbf{d}}$ loads. The synthetic parameters are generated using the standard post-processing steps in (4) with no cyber resilience guarantee. The results reveal that the load redistribution attacks are as effective on synthetic parameters as on the original loads, thus motivating the cyber resilient obfuscation by means of Alg. 1 and 2.
TABLE I: Average outcomes of load redistribution attacks [\$ 1,000]
Testbed Load $C_{\text{opf}}$ $C_{\text{att}}^{\text{BO}}$ (for varying $\eta$)
±$\pm 5\%$$\pm 10\%$$\pm 15\%$
5_pjm actual, $\mathbf{d}$ 88.2 92.5 100.0 108.1
synth., $\tilde{\mathbf{d}}$ 87.4 92.4 100.0 108.1
14_ieee actual, $\mathbf{d}$ 4.80 4.93 5.06 5.19
synth., $\tilde{\mathbf{d}}$ 4.78 4.93 5.03 5.17
24_ieee actual, $\mathbf{d}$ 227.2 255.0 283.0 311.1
synth., $\tilde{\mathbf{d}}$ 212.5 242.3 259.1 274.9
118_ieee actual, $\mathbf{d}$ 237.0 252.4 256.4 259.8
synth., $\tilde{\mathbf{d}}$ 225.1 229.1 238.8 241.0
V-B Insights from the Small PJM 5-Bus Testbed
We test the CRO Alg. 1 in mitigating the attack damage. We generate $1,000$ synthetic loads using the standard post-processing (PP) in (4) and $1,000$ synthetic loads from the CRO assuming $\eta=5\%$. The histograms of the normal and post-attack OPF costs are shown in Fig. 1. Their range becomes wider as load adjacency (and hence the noise) increases. For the standard post-processing (PP) (top row), we observe a notable shift of the post-attack histogram to the right relative to the cost of normal operations, confirming the results from Tab. I. The attacks calibrated on the outcomes of the CRO algorithm result in no extra OPF cost, as the histograms of the normal and post-attack cost overlap (bottom row). Thus, when attacks are calibrated on CRO results, the adversary sees no gain from launching an attack.
Table II shows the impact of the trade-off parameter $\beta$ on the CRO algorithm. The load redistribution attack demonstrate notable damage when disregarding attacks in the CRO algorithm $(\beta=0)$. On the other hand, as long as $\beta$ exceeds the regularization weight $\gamma$, the source grid remains immune to attacks. This trade-off is “flat” as we model the linear OPF costs; we expect it to be smoother for quadratic OPF costs, which is a subject of future investigation.
Figure 1: Histograms of normal and post-attack OPF costs in the PJM 5-bus systems. Blue and red dotted lines represent the average OPF costs on synthetic load parameters in normal and post-attack scenarios, respectively. Top row: histograms resulting from the standard post-processing based on (4). Bottom row: histograms resulting from the CRO algorithm.
TABLE II: OPF costs induced on synthetic load vectors $\tilde{\mathbf{d}}_{\text{cro}}$ for varying trade-off parameters $\beta$ and attack assumptions. Attack magnitudes $\eta=5\%$.
Trade-off Parameters $\alpha=20$MW $\alpha=100$MW $\alpha=200$MW
$C_{\text{opf}}$ $C_{\text{att}}^{\text{BO}}$ $C_{\text{opf}}$ $C_{\text{att}}^{\text{BO}}$ $C_{\text{opf}}$ $C_{\text{att}}^{\text{BO}}$
$\beta\in[0,\gamma)$ 88.2 92.9 87.3 91.5 84.5 88.2
$\beta\in[\gamma,\infty)$ 88.2 88.2 87.3 87.3 84.5 84.5
V-C Large-Scale Applications with CRO-Exp
The post-processing optimization (1) in CRO is difficult to scale to large systems. As shown in Fig. 2, the number of variables and complementarity constraints grow with the size of the testbed. The CRO-Exp Alg. 2 reduces the problem by several orders of magnitude as it only considers a subset of $\tau$ worst-case constraints in the attack. Fig. 3 shows the damage of attacks calibrated on synthetic loads released by CRO-Exp for three large testbeds. The increase of $\tau$ reduces the attack damage. Notably, $\tau=5$ suffices to minimize the attack damage, showing no improvement of cyber resilience beyond this threshold. This is due to the fact that only the attacks on a limited number of constraints can greatly increase the OPF cost. Moreover, the selection of the worst-case constraints in Step 3 of Alg. 2 becomes less informative with more noise, which only increases in $\tau$, as per Theorem 5.
Figure 2: Number of variables and complementarity constraints in CRO and CRO-Exp ($\tau=5$) post-processing across four test systems (log-scale).
Figure 3: Outcomes of the BO load redistribution attack calibrated on synthetic CRO-Exp loads for varying number of the worst-case constraints $\tau$. $\tau=0$ means the synthetic dataset generated by the standard post-processing in (4). Adjacency $\alpha$ are determined in percentages of the average load in the testbed. Attack magnitudes are for $\eta=15\%$ in IEEE 14-bus system and $\eta=5\%$ in IEEE 24-bus and 118-bus systems. Red lines represent the mean value, the blue area represents the 80% confidence interval.
VI Conclusion
We developed algorithms for synthesizing credible grid parameters from real-world systems for OPF analysis. Similar to existing DP algorithms, they obfuscate loads by injecting Laplacian noise and using post-processing; however, they differ in a post-processing stage which optimizes for the trade-off between modeling fidelity (OPF cost consistency) and the resilience of source grids to cyber attacks. Our results reveal that these trade-offs are “flat”, meaning resilience can be achieved with little to no impact on the fidelity of the synthetic data. We also found that the post-processing formulation can be reduced with no loss of resilience using the exponential mechanism to select only important constraints for the attack. Inspired by these observations, future work aims to further investigate these trade-offs in the OPF setting with nonlinear (quadratic) costs and a broader class of attack models amenable to optimization-based representation.
-A Proof of Proposition 3
For any feasible load vectors $\mathbf{d}$, we have
$C_{\text{opf}}^{\text{BO}}(\mathbf{d})$ $=\underset{{\mathbf{x}}}{\text{minimize}}\quad\mathbf{c}^{\top}\mathbf{x}$ (12)
subject to $\mathbf{a}_{k}^{\top}\mathbf{x}+\mathbf{b}_{k}^{\top}(\mathbf{d}+\bm{\gamma}^{\star})+e_{k}\leqslant 0\;:\omega_{k}$
$C_{\text{opf}}^{\text{RO}}(\mathbf{d})$ $=\underset{{\mathbf{x}}}{\text{minimize}}\quad\mathbf{c}^{\top}\mathbf{x}$ (13)
subject to $\mathbf{a}_{k}^{\top}\mathbf{p}+\mathbf{b}_{k}^{\top}(\mathbf{d}+\bm{\delta}_{k}^{\star})+e_{k}\leqslant 0\;:\xi_{k},$
where ${\gamma}^{\star}$ and $\{\bm{\delta}_{1}^{\star},...,\bm{\delta}_{k}^{\star}\}$ are the optimal attack vectors, belonging to the same admissible set $\Delta$. We show the correspondence between the objectives via Lagrange duality of problems (12) and (13). Their dual problems are
$\underset{{\omega_{k}}}{\text{maximize}}\quad$ $\textstyle\sum_{k}\omega_{k}(\mathbf{b}_{k}^{\top}\bm{\gamma}^{\star}+\mathbf{b}_{k}^{\top}\mathbf{d}+e_{k})$ (14a)
subject to $\mathbf{c}+\textstyle\sum_{k}\omega_{k}\mathbf{a}_{k}=\bm{0}\quad\forall k,$ (14b)
$\underset{{\xi_{k}}}{\text{maximize}}\quad$ $\textstyle\sum_{k}\xi_{k}(\mathbf{b}_{k}^{\top}\bm{\delta}_{k}^{\star}+\mathbf{b}_{k}^{\top}\mathbf{d}+e_{k})$ (15a)
subject to $\mathbf{c}+\textstyle\sum_{k}\xi_{k}\mathbf{a}_{k}=\bm{0}\quad\forall k,$ (15b)
respectively. To relate the costs, we thus need to show that
$\textstyle\sum_{k}\xi_{k}(\mathbf{b}_{k}^{\top}\bm{\delta}_{k}^{\star}+\mathbf{b}_{k}^{\top}\mathbf{d}+e_{k})$
$\quad\quad\quad\geqslant\textstyle\sum_{k}\omega_{k}(\mathbf{b}_{k}^{\top}\bm{\gamma}^{\star}+\mathbf{b}_{k}^{\top}\mathbf{d}+e_{k}).$ (16)
From dual constraints (15b) and (14b), we have the equality
$\textstyle\sum_{k}\xi_{k}\mathbf{a}_{k}=\textstyle\sum_{k}\omega_{k}\mathbf{a}_{k},\quad\forall k=1,\dots,K,$ (17)
for which $\xi_{k}=\omega_{k},\forall k=1,\dots,K$ is one possible solution. By Assumption 1, the OPF problem (1) satisfies LICQ, and so do problems (12) and (13) for any attack vector. Per [19, Theorem 2], for any feasible dispatch $\mathbf{p}$, the dual solution to problems (12) and (13) is unique. Hence, $\xi_{k}=\omega_{k}$ is not only one but also unique solution to (17). Since the duals are non-negative in optimality, it suffices to show that
$\mathbf{b}_{k}^{\top}\bm{\delta}_{k}^{\star}\geqslant\mathbf{b}_{k}^{\top}\bm{\gamma}^{\star},\quad\forall k=1,\dots,K,$ (18)
for equation (16) to hold. We recall that the optimal attack vectors $\bm{\delta}_{1}^{\star},\dots,\bm{\delta}_{K}^{\star}$ come from the robust optimization (7), so the left-hand side is $\mathbf{b}_{k}^{\top}\bm{\delta}_{k}^{\star}=\underset{{\bm{\delta}_{k}\in\Delta}}{\text{maximize}}\;\mathbf{b}_{k}^{\top}\bm{\delta}_{k}$. On the other hand, the worst case for this inequality is when $\bm{\gamma}$ maximizes the right-hand side of (18), so we need to show that
$\underset{{\bm{\delta}_{k}\in\Delta}}{\text{maximize}}\;\mathbf{b}_{k}^{\top}\bm{\delta}_{k}\geqslant\underset{{\bm{\gamma}\in\Delta}}{\text{maximize}}\;\mathbf{b}_{k}^{\top}\bm{\gamma},\quad\forall k=1,\dots,K.$ (19)
Since the attack vectors belong to the same set $\Delta$, (19) holds as equality in the worst case. Therefore, inequality (18), and hence (16), always hold, completing the proof.
-B Complete Formulation of the CRO Post Processing
The complete formulation of (1) with the Karush-Kuhn-Tucker conditions (KKTs) of embedded problems is
$\underset{{}}{\text{minimize}}\;$ $\lVert\tilde{C}_{\text{opf}}-\mathbf{c}^{\top}\mathbf{x}_{1}\rVert_{1}+\beta\lVert\tilde{C}_{\text{opf}}-\mathbf{c}^{\top}\mathbf{x}_{2}\rVert_{1}+\gamma\lVert\tilde{\mathbf{d}}-\tilde{\mathbf{d}}^{0}\rVert_{1}$
subject to
KKTs of RO approxiamtion (8) $\left\{\begin{array}[]{ll}\overline{\bm{\mu}}_{k}^{\top}\overline{\bm{\delta}}-\underline{\bm{\mu}}_{k}^{\top}\underline{\bm{\delta}}\leqslant-\mathbf{a}_{k}^{\top}\mathbf{x}_{1}-\mathbf{b}_{k}^{\top}\tilde{\mathbf{d}}-e_{k}\\ \mathbf{b}_{k}-\overline{\bm{\mu}}_{k}+\underline{\bm{\mu}}_{k}-\bm{1}\lambda_{k}=\bm{0}\\ \underline{\bm{\mu}}_{k},\overline{\bm{\mu}}_{k}\geqslant\bm{0}\\ \mathbf{c}+\sum_{k}\theta_{k}\mathbf{a}_{k}=\mathbf{0}\\ \theta_{k}\overline{\bm{\delta}}-\bm{\zeta}_{k}-\overline{\bm{\pi}}_{k}=\bm{0}\\ \theta_{k}\underline{\bm{\delta}}-\bm{\zeta}_{k}+\underline{\bm{\pi}}_{k}=\bm{0}\\ \bm{1}^{\top}\bm{\zeta}_{k}=0\\ 0\leqslant\theta_{k}\;\perp(-\overline{\bm{\mu}}_{k}^{\top}\overline{\bm{\delta}}+\underline{\bm{\mu}}_{k}^{\top}\underline{\bm{\delta}}-\mathbf{a}_{k}^{\top}\mathbf{x}_{1}\\ \quad\quad\quad\quad\quad\!-\mathbf{b}_{k}^{\top}\tilde{\mathbf{d}}-e_{k})\geqslant 0\\ \bm{0}\leqslant\underline{\pi}_{k}\;\perp\underline{\bm{\mu}}_{k}\geqslant\bm{0}\\ \bm{0}\leqslant\overline{\pi}_{k}\;\perp\overline{\bm{\mu}}_{k}\geqslant\bm{0}\\ \theta_{k},\overline{\bm{\pi}}_{k},\underline{\bm{\pi}}_{k},\bm{\zeta}\geqslant\bm{0}\end{array}\right\}$
KKTs of OPF (2) $\left\{\begin{array}[]{ll}\mathbf{a}_{k}^{\top}\mathbf{x}_{2}+\mathbf{b}_{k}^{\top}\tilde{\mathbf{d}}+e_{k}\leqslant 0\\ \mathbf{c}+\sum_{k}\nu_{k}\mathbf{a}_{k}=0\\ 0\leqslant\nu_{k}\;\perp(-\mathbf{a}_{k}^{\top}\mathbf{x}_{2}-\mathbf{b}_{k}^{\top}\tilde{\mathbf{d}}-e_{k})\geqslant 0\\ \nu_{k}\geqslant 0\end{array}\right\}$
$\forall k=1,\dots,K$, where $\mathbf{c}^{\top}\mathbf{x}_{1}$ and $\mathbf{c}^{\top}\mathbf{x}_{2}$ represents the OPF cost in post-attack and normal conditions, respectively. Here, the $\perp$ denotes complementarity conditions.
-C Proof of Theorem 4
CRO uses the real data in the following computations:
  1. 1.
    Step 1 adds Laplacian noise with magnitude $\alpha/\varepsilon_{2}$ to an identity query, whose sensitivity is $\alpha$. By the sequential composition rule [4], this computation is $\varepsilon_{1}-$DP.
  2. 2.
    Step 2 adds Laplacian noise with parameter $(\alpha\overline{c})/\varepsilon_{2}$. Since the sensitivity of OPF cost is $\alpha\overline{c}$ as shown in Section II.B, this computation is $\varepsilon_{2}-$DP.
Since the post-processing optimization in Step 3 only uses obfuscated data, it will not induce any privacy loss due to post-processing immunity [4]. Per the sequential composition rule, the total privacy loss of the algorithm is $\varepsilon_{1}+\varepsilon_{2}$, which adds up to $\varepsilon$ when we take $\varepsilon_{1}=\varepsilon/2,\varepsilon_{2}=\varepsilon/2$.
-D Proof of Theorem 5
The algorithm queries data in the following computations:
  1. 1.
    Following the similar arguments from Appendix C, Step 1 is $\varepsilon_{1}-$DP and Step 2 is $\varepsilon_{2}-$DP
  2. 2.
    The worst-case constraints are estimated using $\tau$ iterations of the report-noisy-max algorithm in Step 3; each iteration injects the Laplacian noise with magnitude $\alpha\overline{c}$ providing $\varepsilon_{3}-$DP, and the whole report-noisy-max algorithm is $\tau\varepsilon_{3}-$DP.
As the post-processing optimization in Step 4 only uses obfuscated numerical data $\tilde{\mathbf{d}}^{0},\tilde{C}_{\text{opf}}$ and non-numerical data $\mathcal{K}$, it is immune to privacy loss. The accumulated privacy loss of Alg. 2 is $\varepsilon_{1}+\varepsilon_{2}+\tau\varepsilon_{3}$, which amounts to $\varepsilon$ when setting $\varepsilon_{1}=\varepsilon/3,\varepsilon_{2}=\varepsilon/3$ and $\varepsilon_{3}=\varepsilon/(3\tau)$.